OpenStudy (anonymous):

Put the equation in the form ax2 + bx + c = 0. 3x2+4x=(3x-1)2-2x2

OpenStudy (cwrw238):

3x2 + 4x = 9x^2 + 1 -6x - 2x^2 can you continue from here?

OpenStudy (anonymous):

its what comes after that gets confusing

OpenStudy (cwrw238):

well first bring all terms to right of '=': 0 = 9x^2 - 2x^2 - 3x^2 - 4x - 6x + 1 0 = 4x^2 - 10x + 1 or 4x^2 - 10x + 1 = 0

OpenStudy (cwrw238):

- bring all the like terms together and simplify

OpenStudy (anonymous):

ok...im working it out on paper as well nd i got that

OpenStudy (cwrw238):

good job

OpenStudy (anonymous):

thanks a lot for help

OpenStudy (cwrw238):

no probs